OpenStudy (anonymous):

Find the value of C. Express the answer as a simplified radical like 5sqrt(2).

OpenStudy (anonymous):

i got 725^2 but that was wrong

OpenStudy (anonymous):

wait is that right???!!!

hartnn (hartnn):

but, whats the expression for C ? i mean there's nothing to simplify.....

OpenStudy (anonymous):

there you go :p

OpenStudy (anonymous):

so is it 5^29?

hartnn (hartnn):

use pythagoras theorem, \(C^2 =25^2+10^2=...?\)

hartnn (hartnn):

if you meant \(\huge 5\sqrt {29}\) then you are correct!

OpenStudy (anonymous):

ya and the square root of 725 is 5^29 so that is correct then the answer is 5^29?

hartnn (hartnn):

yes, the answer is \(5\sqrt{299}\)

hartnn (hartnn):

***\(5\sqrt{29}\)

hartnn (hartnn):

oh, and hey, \(\Huge \mathcal{\text{Welcome To OpenStudy}\ddot\smile} \)

OpenStudy (anonymous):

lol thanks your awesome! xD

hartnn (hartnn):

haha, you too ;)

OpenStudy (anonymous):

LOL I know im awesome jk =p...hey how do you make the square root sign

hartnn (hartnn):

by typing \sqrt here, the code `\( \sqrt 2 \)` will be displayed as \( \sqrt 2 \)

hartnn (hartnn):

or you can use the " \(\sum \)Equation " Tab below to write square root

hartnn (hartnn):

in there 4th row, 1st column has \(\sqrt .\) sign

OpenStudy (anonymous):

lol ok well find me... and ill delete that info now :D